OpenStudy (lovelyharmonics):

How many atoms of cobalt would there be in a 2.2-gram sample of cobalt? 4.64 x 1021 atoms 2.25 x 1022 atoms 1.61 x 1025 atoms 7.81 x 1025 atoms @ganeshie8

ganeshie8 (ganeshie8):

http://www.chemicool.com/elements/cobalt.html

OpenStudy (lovelyharmonics):

wouldnt it be b?

ganeshie8 (ganeshie8):

atomic weight of cobalt \(\approx \) 59

ganeshie8 (ganeshie8):

that means there are \(\large 6.022 \times 10^{23}\) atoms in \(\large 59 ~grams\) of cobals

ganeshie8 (ganeshie8):

\[\large \dfrac{6.022\times 10^{23}}{59} = \dfrac{x}{2.2}\]

OpenStudy (lovelyharmonics):

2.2 grams/(58.9g/mole)=0.0374mole 6.023 x 10^ 23 atoms / mole 0.0374 moles (6.023 x 10^23 atoms/mole)=2.25 x 10^22 atoms of cobalt.

ganeshie8 (ganeshie8):

\[\large x = \dfrac{6.022\times 10^{23}}{59}\times 2.2 \]

ganeshie8 (ganeshie8):

b is right !